# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ 8.6 Perturbed Dirac operators and the Cauchy singular integral operator
Then $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{2}(\mathbb{R}),D_{\gamma})$ is a locally compact Banach spectral triple over the algebra $\mathrm{C}_{0}^{\infty}(\mathbb{R})$.
@@ 8.7 Banach Fredholm modules arising from pseudo-differential operators
Proof : By [BaG82b, Proposition 2.4, p. 143] and its proof, the operator $\operatorname{\mathrm{sign}}D\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$ induces a pseudo-differential operator $\operatorname{\mathrm{sign}}P\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ of order 0. Recall that a multiplication operator $M_{f}\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ is a differential operator of order 0, hence a pseudo-differential operator of order 0. It is easy to check5554. We warn the reader that the commutator of two pseudo-differential operators on a vector bundle $E$ of order $m$ and $m^{\prime}$ is not necessarily of order $m+m^{\prime}-1$. using the formula [Tre80, (7.2) p. 75] that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ is a pseudo-differential operator of order 1-1. Since it has negative order, it induces by Lemma 8.36 a compact operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, it equally induces by Theorem 8.37 a bounded operator on the Banach space $\mathrm{L}^{q}(M,E)$ for any $1<q<\infty$. By interpolation [KZPP76, Theorem 3.11 p. 58] relying on the formula (8.45), we conclude that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ induces a compact operator on the Banach space $\mathrm{L}^{p}(M,E)$. By approximation, we see that it is true more generally for any function $f\in\mathrm{C}(M)$.
@@ 9.4 Banach Fredholm modules on reduced group $\mathrm{C}^{*}$-algebras
If an orientable, irreducible6665. A 3-manifold $M$ is irreducible if every smooth 2-sphere in $M$ bounds a 3-dimensional ball in $M$. 3-manifold $M$ has infinite first homology group, then by [ClR16, Theorem 5.3 p. 66] its fundamental group $\pi_{1}(M)$ is locally indicable, hence left-orderable.
@@
Let $\Sigma^{b}_{g,n}$ be an orientable connected surface with genus $g$, $b$ boundary components, and $n$ punctures7776. Recall that $P=\{p_{1},\ldots,p_{n}\}$ is a finite subset of distinct points ¡¡removed¿¿ from the surface. $p_{1},\ldots,p_{n}$. If $b>0$ and if the Euler characteristic of $\Sigma^{b}_{g,n}$ satisfies $\chi(\Sigma^{b}_{g,n})<0$ then by [ClR16, Theorem 7.29 p. 107], the mapping class group $\mathrm{Mod}(\Sigma^{b}_{g,n})$ is left-orderable. This group is the group of isotopy classes of orientation-preserving homeomorphisms that restrict to the identity on the boundary of $\Sigma^{b}_{g,n}$.
@@
Moreover, we can consider the subalgebra $\mathcal{D}\overset{\mathrm{def}}{=}\mathbb{C}1$ of the group von Neumann algebra $\mathrm{VN}(G)$. The linear map $\mathrm{E}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$, $f\mapsto\tau(f)1$ is a trace preserving normal faithful conditional expectation onto the ¡¡diagonal¿¿ subalgebra $\mathcal{D}$. If $G$ is left-orderable and endowed with a left-order $\preceq$ then it is essentially8887. The proposed definition in $\mathrm{H}^{\infty}(\mathrm{VN}(G))$ is slightly incorrect since the ¡¡noncommutative Fourier series¿¿ of elements in $\mathrm{VN}(G)$ does not converges in general. observed in [GPX22] that
@@ Acknowledgment
The author would like to thank Hermann Schulz-Baldes for bringing his recent book [ScS22] to our attention on his own initiative.

### v4

@@ 8.6 Perturbed Dirac operators and the Cauchy singular integral operator
Then $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{2}(\mathbb{R}),D_{\gamma})$ is a possibly kernel-degenerate locally compact Banach spectral triple.
@@ 8.7 Banach Fredholm modules arising from pseudo-differential operators
Proof : By [BaG82b, Proposition 2.4, p. 143] and its proof, the operator $\operatorname{\mathrm{sign}}D\colon\mathrm{L}^{2}(M,E)\to\mathrm{L}^{2}(M,E)$ induces a pseudo-differential operator $\operatorname{\mathrm{sign}}P\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ of order 0. Recall that a multiplication operator $M_{f}\colon\mathrm{C}^{\infty}(M)\to\mathrm{C}^{\infty}(M)$ is a differential operator of order 0, hence a pseudo-differential operator of order 0. It is easy to check6665. We warn the reader that the commutator of two pseudo-differential operators on a vector bundle $E$ of order $m$ and $m^{\prime}$ is not necessarily of order $m+m^{\prime}-1$. using the formula [Tre80, (7.2) p. 75] that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ is a pseudo-differential operator of order 1-1. Since it has negative order, it induces by Lemma 8.48 a compact operator on the Hilbert space $\mathrm{L}^{2}(M,E)$. Moreover, it equally induces by Theorem 8.49 a bounded operator on the Banach space $\mathrm{L}^{q}(M,E)$ for any $1<q<\infty$. By interpolation [KZPP76, Theorem 3.11 p. 58] relying on the formula (8.77), we conclude that the commutator $\operatorname{\mathrm{sign}}P,M_{f}]$ induces a compact operator on the Banach space $\mathrm{L}^{p}(M,E)$. By approximation, we see that it is true more generally for any function $f\in\mathrm{C}(M)$.
@@ 9.4 Banach Fredholm modules on reduced group $\mathrm{C}^{*}$-algebras
If an orientable, irreducible7776. A 3-manifold $M$ is irreducible if every smooth 2-sphere in $M$ bounds a 3-dimensional ball in $M$. 3-manifold $M$ has infinite first homology group, then by [ClR16, Theorem 5.3 p. 66] its fundamental group $\pi_{1}(M)$ is locally indicable, hence left-orderable.
@@
Let $\Sigma^{b}_{g,n}$ be an orientable connected surface with genus $g$, $b$ boundary components, and $n$ punctures8887. Recall that $P=\{p_{1},\ldots,p_{n}\}$ is a finite subset of distinct points ¡¡removed¿¿ from the surface. $p_{1},\ldots,p_{n}$. If $b>0$ and if the Euler characteristic of $\Sigma^{b}_{g,n}$ satisfies $\chi(\Sigma^{b}_{g,n})<0$ then by [ClR16, Theorem 7.29 p. 107], the mapping class group $\mathrm{Mod}(\Sigma^{b}_{g,n})$ is left-orderable. This group is the group of isotopy classes of orientation-preserving homeomorphisms that restrict to the identity on the boundary of $\Sigma^{b}_{g,n}$.
@@
Moreover, we can consider the subalgebra $\mathcal{D}\overset{\mathrm{def}}{=}\mathbb{C}1$ of the group von Neumann algebra $\mathrm{VN}(G)$. The linear map $\mathrm{E}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$, $f\mapsto\tau(f)1$ is a trace preserving normal faithful conditional expectation onto the ¡¡diagonal¿¿ subalgebra $\mathcal{D}$. If $G$ is left-orderable and endowed with a left-order $\preceq$ then it is essentially9998. The proposed definition in $\mathrm{H}^{\infty}(\mathrm{VN}(G))$ is slightly incorrect since the ¡¡noncommutative Fourier series¿¿ of elements in $\mathrm{VN}(G)$ does not converges in general. observed in [GPX22] that
@@ Acknowledgment
The author would like to thank Hermann Schulz-Baldes for bringing his recent book [ScS22] to our attention on his own initiative and Anton Savin and Markus Haase for short discussions.